Makala2fine:

Solve given system 4x-5y=40 x=3y+24

Makala2fine:

@xxaikoxx

Luigi0210:

One way you could do it is by substitution. Subbing in \(\color{red}{x}=3y+24\) for \( 4\color{red}{x} -5y=40 \) Giving you something like this: \(\large 4(3y+24)-5y =40 \) Solve for y: \(\large 12y+96-5y=40 \) \(\large 7y=56 \) \(\large y= 8\) Now that you have \(\large \color{gold}{y}\), you can plug that back in and solve for \(\large \color{red}{x}\).

gdubb08:

The only thing I got outta this is when I looked at x, I wondered y.

Makala2fine:

@luigi0210 wrote:
One way you could do it is by substitution. Subbing in \(\color{red}{x}=3y+24\) for \( 4\color{red}{x} -5y=40 \) Giving you something like this: \(\large 4(3y+24)-5y =40 \) Solve for y: \(\large 12y+96-5y=40 \) \(\large 7y=56 \) \(\large y= 8\) Now that you have \(\large \color{gold}{y}\), you can plug that back in and solve for \(\large \color{red}{x}\).
Thank you

surjithayer:

4x-5y=40 x=3y+24 4(3y+24)-5y=40 12y+96-5y=40 7y=40-96 7y=-56 y=-56/7=-8 x=3(-8)+24 x=-24+24 x=0 x=0,y=-8